If a ball is thrown into the air with a velocity of 40 ft/s, its height (in feet) after t seconds is given by y = 40t − 16t2. Find the velocity when t = 2 seconds.


1
Expert's answer
2021-10-11T15:14:55-0400

Height of the function is given by "y=40t-16t^2"

Require to find the velocity when "t=2" seconds


To find the velocity, find "\frac{dy}{dt}" from the given function "y=40t-16t^2"

Now "y=40t-16t^2\Rightarrow \frac{dy}{dt}=\frac{d}{dt}\left [40t-16t^2 \right ]"

"\Rightarrow \frac{dy}{dt}= 40(1)-16(2t)"

"\Rightarrow \frac{dy}{dt}= 40-32t"


So, velocity is given by "v(t)=40-32t"

Now let us find the velocity when "t=2" seconds


"t=2\Rightarrow v(2)=40-32(2)=40-64=-24"


Therefore, "v(2)=-24 ft\/s"